Simplify X^p+(1/2)q × Y^2p-q ÷ (XY^2)^p × √X^q
1
Expert's answer
2020-07-07T21:07:49-0400

"\dfrac{x^{p+\frac{1}{2}q} \cdot y^{(2p-q)}}{(x\cdot y^2)^p \cdot \sqrt{x}^q} =
\dfrac{x^p\cdot x^{\frac{1}{2}q} \cdot y^{2p}\cdot y^{-q}}{x^p\cdot y^{2p} \cdot x^{\frac{1}{2}q}} =
y^{-q}"